Suppose the risk-free rate is 1.35% and an analyst assumes a market risk premium of 6.73%. Firm A just paid a dividend of $1.24 per share. The analyst estimates the β of Firm A to be 1.21 and estimates the dividend growth rate to be 4.87% forever. Firm A has 293.00 million shares outstanding. Firm B just paid a dividend of $1.89 per share. The analyst estimates the β of Firm B to be 0.77 and believes that dividends will grow at 2.18% forever. Firm B has 186.00 million shares outstanding. What is the value of Firm A?

With calculator steps
Answer #1

Required rate of return as per CAPM = Risk free rate + beta*Market risk premium

Required rate of return on Firm A = 1.35% + 1.21*6.73% = 9.4933%

Value of share = Expected Dividend/(Required return - growth rate)

= 1.24(1+4.87%)/(9.4933% - 4.87%)

= $28.1268

Hence, value of firm A = 28.1268*293 million

= $8,241.16 million